Calculate the change of Gibbs energy when 3.0 mol H2 mixes with 1.0 mol N2 at the same pressure with the volume of the vessel adjusted accordingly.
1
Expert's answer
2020-11-04T14:15:17-0500

Solution.

"\chi (H2) = \frac{3}{3+1} = 0.75"

"\chi (N2) = \frac{1}{3+1} = 0.25"

n = 4 mol

"\Delta G = nRT(\chi(H2) \times ln(\chi(H2)) + \chi(N2) \times ln(\chi(N2)))"

"\Delta G = 4 \times 8.31 \times 298(0.75 \times ln(0.75) + 0.25 \times ln(0.25)) = -5570.22 J/mol"

Answer:

"\Delta G = -5570.22 J/mol"